# Model-Free Assessment of Simulator Fidelity via Quantile Curves

### v1

@@ Appendix A Technical Lemmas
B.1 Proof to Theorem 3.1:
As a setup, we work conditionally on $\mathcal{G}_{j}:=\sigma(\psi_{j},\hat{q}_{j},n_{j}),\mathcal{G}:=\sigma\big(\{\mathcal{G}_{j}\}_{j=1}^{m}\big).$ Here $\mathcal{G}_{j}$ is the $\sigma$-field collecting all “human-side” information for scenario $j$. Thus $\mathcal{G}:=\sigma(\{\mathcal{G}_{j}\}_{j=1}^{m})$ represents the joint information from all scenarios that we treat as fixed when analyzing the remaining randomness coming from the simulator estimates $\hat{p}_{j}$. For brevity, we abuse notation and write $p(\psi_{j}),q(\psi_{j})$ as $p_{j},q_{j}$, and similarly $\hat{p},\hat{q}$. We also ignore the superscript $k$ to simplify the proof, and discussions on the dependence on $k$ can be found in Section 3. In addition, for any quantities $\{\Delta_{j}\}_{j=1}^{m}$, we denote the sorted version as $\{\Delta_{(i)}\}_{i=1}^{m}$, ie. $\Delta_{(1)}\leq\cdots\leq\Delta_{(m)}.$ For any sequence $\{\Delta_{j}\}_{j=1}^{m}$, let $\Delta_{(1)}\leq\cdots\leq\Delta_{(m)}$ denote its order statistics.
@@
We first consider a fixed level of confidence $\bar{\alpha}$, then extend to $\bar{\alpha}$ holding uniformly across $(0,1)$. Throughout the proof, $\alpha\in(0,1)$ denotes a generic quantile index (a function argument) and is distinct from the target coverage level $\bar{\alpha}$.
@@
However, for the above argument to hold, we would need to know $p_{j}$, instead we only have the estimated $\hat{p}_{j}$. Therefore, we instead construct a randomized pseudo-gaps. Specifically, we set a level of coverage $\gamma=\tfrac{1}{2}$, and define the two gap terms as
@@
where $\mathcal{C}_{j}\!\left(\hat{p}_{j},\gamma\right)\subset\Theta$ are data-driven confidence sets satisfying $\mathbb{P}(p_{j}\in\mathcal{C}_{j}\!\left(\hat{p}_{j},\gamma\right)\big|\psi_{j},n_{j})\geq\tfrac{1}{2}.$
By Assumption 3.2 and the compactness of the confidence set $C_{j}(\hat{p}_{j},\gamma)\subset\Theta$, Berge’s maximum theorem guarantees that the supremum in the definition of $\hat{\Delta}_{j}$ is attained, hence $\hat{\Delta}_{j}$ is well defined. Therefore, by the coverage property of $C_{j}$, we have
$\mathbb{P}(\hat{\Delta}_{j}\geq{\Delta}_{j}|\mathcal{G}_{j}\cup\sigma(\gamma))\geq\gamma,$
and by $\gamma\perp\!\!\!\perp\mathcal{G}_{j}$ we get $\forall j$
$\mathbb{P}(\hat{\Delta}_{j}\geq{\Delta}_{j}|\mathcal{G}_{j})$ $=\mathbb{E}[\mathbb{P}(\hat{\Delta}_{j}\geq{\Delta}_{j}|\mathcal{G}_{j}\cup\sigma(\gamma)|\mathcal{G}_{j}]$
$\geq\frac{1}{2}.$ (B.2)
@@
$\mathbb{P}\big(\hat{\Delta}_{j}\geq\Delta_{j}\big)=\mathbb{\mathbb{E}}\big[\mathbb{P}(\hat{\Delta}_{j}\geq\Delta_{j}\big|\mathcal{G}_{j})\big]\geq\tfrac{1}{2}.$
Moreover, by the above discussion, conditional on $\mathcal{G}$, the indicators $Y_{j}=\mathbf{1}\{\hat{\Delta}_{j}\geq\Delta_{j}\}$ are independent across $j$.
We will use $\{\hat{\Delta}_{j}\}_{j=1}^{m}$ to create an upper bound of $\bar{V}_{m}(1-\alpha)$, which along side (B.1) will give us our desired envelope. Intuitively, we want to find a larger quantile of $\hat{\Delta}_{j}$ and with (B.1), we can claim an upper bound of $\bar{V}_{m}(1-\alpha)$ with high probability.
@@
We next define $Y_{j}=\mathbb{I}(\hat{\Delta}_{j}\geq\Delta_{j})$. By (B.1): $\mathbb{P}(Y_{j}=1|\mathcal{\mathcal{G}})\geq\frac{1}{2}$. Next, define $T_{\alpha}=\{j\in S_{\alpha}:\hat{\Delta}_{j}\geq\Delta_{j}\}=\{j\in S_{\alpha}:Y_{j}=1\}=\{j:\hat{\Delta}_{j}\geq\Delta_{j}\geq\bar{V}_{m}(1-\alpha)\}$. First, we calculate $\mathbb{E}[T_{\alpha}]$:
$\mathbb{E}[T_{\alpha}|\mathcal{G}]=\mathbb{E}\Big[\sum_{j\in S_{\alpha}}Y_{j}|\mathcal{G}\Big]=\sum_{j\in S_{\alpha}}\mathbb{P}(Y_{j}|\mathcal{G})\geq\frac{1}{2}s.$
Lemma A.2 implies that for any $\delta\in[0,\frac{1}{2}s]$, we have:
$\mathbb{P}\Big(|T_{\alpha}|\leq\frac{1}{2}s-t|\mathcal{G}\Big)$ $\leq\mathbb{P}\Big(|T_{\alpha}|\leq\mathbb{E}[T_{\alpha}]-t|\mathcal{G}\Big)$
@@
where we applied $\mathbb{E}[T_{\alpha}|\mathcal{G}]\geq\tfrac{1}{2}s$ in the first inequality. By setting $t=c\sqrt{s}$:
$\mathbb{P}\Big(|T_{\alpha}|\leq\frac{1}{2}s-t|\mathcal{G}\Big)\leq\exp(-2c^{2})$ (B.3)
@@
$|U_{\alpha}|\geq|T_{\alpha}|\geq\frac{1}{2}s-c\sqrt{s},$
which implies at least $\frac{1}{2}s-c\sqrt{s}$ of the $\hat{\Delta}_{j}$’s are larger than $\bar{V}_{m}(1-\alpha)$ with high probability.
We now analyze what coverage guarantee can we get for the inner probability via order statistics for any $\alpha$. Set $N:=\big\lfloor\tfrac{1}{2}s-c\sqrt{s}\big\rfloor$ and define $\hat{V}_{m}(\alpha):=\inf\{t:\hat{F}_{m}(t)\geq\alpha\}=\Delta_{(\lceil m\alpha\rceil)}$, where $\hat{F}(x)=\frac{1}{m}\sum_{j=1}^{m}\mathbb{I}(\hat{\Delta}_{j}\leq x)$. If at least $N$ sample values exceed $\bar{V}_{m}(1-\alpha)$, then by order-statistics calculus (Lemma A.7)
@@
$\tfrac{1}{2}m\alpha_{\mathrm{eff}}-c\sqrt{m\alpha_{\mathrm{eff}}}-1\ \geq\ m\alpha.$
Define $\alpha_{\mathrm{eff}}(\alpha,c,m):=\inf\{x\in(0,1):\tfrac{1}{2}x-c\sqrt{\tfrac{x}{m}}-\tfrac{1}{m}\geq\alpha\}$. Writing $y=\sqrt{x}$, this is equivalent to $y^{2}-\tfrac{2c}{\sqrt{m}}\,y-\big(\tfrac{2}{m}+2\alpha\big)\geq 0$, so the minimal admissible $y$ is
$y^{*}=\frac{2c/\sqrt{m}+\sqrt{4c^{2}/m+8\alpha+8/m}}{2},\qquad$
@@
$\text{then}\qquad\mathbb{P}(\mathcal{E}_{\alpha})\ \geq\ 1-e^{-2c^{2}}.$
Apply (B.1) at level $1-\alpha_{\mathrm{eff}}(\bar{\alpha},c,m)$:
$\mathbb{P}_{\psi\sim\Psi}\Big(L(p(\psi),\hat{q}(\psi))\leq\bar{V}_{m}(1-\alpha_{\mathrm{eff}})\Big|\mathcal{D}\Big)$
$\geq\ 1-\alpha_{\mathrm{eff}}(\bar{\alpha},c,m)\ -\ \varepsilon_{m}(\delta).$
@@
$\mathbb{P}_{\psi\sim\Psi}\big(L(p(\psi),\hat{q}(\psi))\leq\hat{V}_{m}(1-\bar{\alpha})\,\big|\,\mathcal{D}\big)$
$\geq\ 1-\alpha_{\mathrm{eff}}(\bar{\alpha},c,m)\ -\ \varepsilon_{m}(\delta),$
@@
$\alpha_{\mathrm{eff}}(\bar{\alpha},$ $c,m)=\frac{\big(\tfrac{2c}{\sqrt{m}}+\sqrt{\tfrac{4c^{2}}{m}+8\bar{\alpha}+\tfrac{8}{m}}\big)^{2}}{4}=$
$2\bar{\alpha}+\frac{c}{\sqrt{m}}\sqrt{8\bar{\alpha}+\tfrac{4c^{2}+8}{m}}+\frac{2c^{2}+2}{m}.$
Therefore, as $m\to\infty$, $\alpha_{\mathrm{eff}}(\alpha,c,m)=2\bar{\alpha}+c\sqrt{8\bar{\alpha}/m}+O(m^{-1})$.
@@
By DKW, $\Pr(\mathcal{E}_{\mathrm{DKW}})\geq 1-\delta$, and by the fixed-level argument, $\Pr(\mathcal{E}_{r})\geq 1-e^{-2c^{2}}$ for each $r$. Hence, by a union bound,
$\Pr\Big(\mathcal{E}_{\mathrm{DKW}}\cap\bigcap_{r=1}^{m}\mathcal{E}_{r}\Big)\ \geq\ 1-\delta-me^{-2c^{2}}.$
@@
Applying (3.2) at level $\alpha_{+}$ yields
@@
$\geq 1-2\alpha_{+}-\frac{c}{\sqrt{m}}\sqrt{8\alpha_{+}+\frac{4c^{2}+8}{m}}-\frac{2c^{2}+2}{m}-\varepsilon_{m}(\delta).$
@@
$2\alpha_{+}\leq 2\alpha+\tfrac{2}{m},\qquad\sqrt{8\alpha_{+}+\tfrac{4c^{2}+8}{m}}\leq\sqrt{8\alpha+\tfrac{4c^{2}+16}{m}}.$
Therefore, with probability at least $1-\delta-me^{-2c^{2}}$ over $\mathcal{D}$, the guarantee (3.2) (at $\alpha$ replaced by $\alpha_{+}$) holds uniformly for all $\alpha\in(0,1)$. The form in the main theorem is a simplification with respect to $c$ and $\delta$.
B.2 Corollary to Theorem 3.1:

### v2

@@ Appendix A Technical Lemmas
B.1 Proof for Theorem 3.1
We condition on $\mathcal{G}_{j}:=\sigma(\psi_{j},\hat{q}_{j},n_{j})$ and $\mathcal{G}:=\sigma(\{\mathcal{G}_{j}\}_{j=1}^{m})$. By Assumption 3.1, conditional on $\mathcal{G}$ the additional randomness used to construct $\hat{p}_{j}$ (hence $\hat{\Delta}_{j}$ and the indicators below) is independent across $j$. For brevity, we abuse notation and write $p(\psi_{j}),q(\psi_{j})$ as $p_{j},q_{j}$, and similarly $\hat{p},\hat{q}$. We also ignore the superscript $k$ to simplify the proof, and discussions on the dependence on $k$ can be found in Section 3. Split $\delta=\delta_{\mathrm{DKW}}+\delta_{\mathrm{al}}+\delta_{\mathrm{cnt}}$ with $\delta_{\mathrm{DKW}}=\delta_{\mathrm{al}}=\delta_{\mathrm{cnt}}=\delta/3$, and union bound the three events defined below.
We first show an idealized case where we have access to the oracle $p_{j}$, and hence can obtain the actual sim-to-real discrepancy $\Delta_{j}$. Under this setting, we can directly define $F(t):=\mathbb{P}_{\psi\sim\Psi}(\Delta(\psi)\leq t)$ and
$\bar{F}_{m}(t)$ $:=\frac{1}{m}\sum_{j=1}^{m}\mathbf{1}\{\Delta_{j}\leq t\},$
$\bar{V}_{m}(u)$ $:=\inf\{t:\bar{F}_{m}(t)\geq u\}.$
By the DKW inequality (Lemma A.1), the following event
$\mathcal{E}_{\mathrm{DKW}}:=\Big\{\sup_{t}|\bar{F}_{m}(t)-F(t)|\leq\varepsilon_{m}(\delta)\Big\},$
occur with probability at least $1-\delta_{\mathrm{DKW}}$, where $\varepsilon_{m}(\delta)=\sqrt{\frac{\log(6/\delta)}{2m}}$. On $\mathcal{E}_{\mathrm{DKW}}$, for every $u\in(0,1)$,
$\mathbb{P}_{\psi\sim\Psi}\big(\Delta(\psi)\leq\bar{V}_{m}(u)\ \big|\ \mathcal{D}\big)\geq u-\varepsilon_{m}(\delta).$ (B.1)
Now in our dataset, we do not have access to the above oracle $\Delta_{j}$. Therefore, we control for $\Delta_{j}$ using the pseudo-discrepancy $\hat{\Delta}_{j}$. Define $Y_{j}:=\mathbf{1}\{\hat{\Delta}_{j}\geq\Delta_{j}\}.$ By the coverage of $\mathcal{C}_{j}(\hat{p}_{j},\gamma_{j})$ and the definition of $\hat{\Delta}_{j}$,
$\mathbb{P}(Y_{j}=1\mid\mathcal{G}_{j})\geq\gamma_{j}.$
For $r=1,\dots,m$, let $\alpha_{r}:=r/m$ and define the oracle right-tail set
$S_{r}:=\{j\in[m]:\Delta_{j}\geq\bar{V}_{m}(1-\alpha_{r})\},\quad s_{r}:=|S_{r}|.$
By definition of empirical quantiles, $s_{r}\geq r$. We then introduce link from the oracle set to the pseudo-discrepancies by defining counts
$T_{r}:=\sum_{j\in S_{r}}Y_{j},\quad U_{r}:=\bigl|\{j:\hat{\Delta}_{j}\geq\bar{V}_{m}(1-\alpha_{r})\}\bigr|.$
By construction, $U_{r}\geq T_{r}$. Notice that $U_{r}$ will give us the amount of $\hat{\Delta}$ required to achieve the tail count of the empirical oracle quantile function, hence we will try to create a lower bound for $U_{r}$ via $T_{r}$.
Under Assumption 3.1, $\{n_{j}\}$ are detereministic, and hence independent of $\{\psi_{j}\}$ and therefore independent of $\{\Delta_{j}\}$; consequently, the random index set $S_{r}$ (which is measurable w.r.t. $\{\Delta_{j}\}_{j=1}^{m}$) is independent of the vector $(\gamma_{1},\dots,\gamma_{m})$. Conditional on the realized multiset $\{\gamma_{j}\}_{j=1}^{m}$, the subcollection $\{\gamma_{j}:j\in S_{r}\}$ is thus a simple random sample without replacement from $\{\gamma_{j}\}_{j=1}^{m}$. By Hoeffding–Serfling Bardenet and Maillard (2015), for each $r$ and each $\tau>0$,
$\mathbb{P}\Big(\frac{1}{s_{r}}\sum_{j\in S_{r}}\gamma_{j}\leq\bar{\gamma}_{m}-\tau\Big)\leq\exp(-2s_{r}\tau^{2}).$
Set
$\tau_{r}:=\sqrt{\frac{\log(m/\delta_{\mathrm{al}})}{2s_{r}}}\leq\sqrt{\frac{\log(3m/\delta)}{2r}}=:\varepsilon_{\mathrm{al}}(\alpha_{r},\delta),$
and union bound over $r=1,\dots,m$ to obtain an event $\mathcal{E}_{\mathrm{al}}$ with $\mathbb{P}(\mathcal{E}_{\mathrm{al}})\geq 1-\delta_{\mathrm{al}}$ on which, for all $r$,
$\sum_{j\in S_{r}}\gamma_{j}\ \geq\ s_{r}\big(\bar{\gamma}_{m}-\varepsilon_{\mathrm{al}}(\alpha_{r},\delta)\big).$ (B.2)
Conditional on $\mathcal{G}$, the indicators $\{Y_{j}\}_{j=1}^{m}$ are independent and bounded in $[0,1]$, so Lemma A.2 implies that for any $t>0$,
$\mathbb{P}\Big(T_{r}\leq\mathbb{E}[T_{r}\mid\mathcal{G}]-t\Big|\mathcal{G}\Big)\leq\exp\Big(-\frac{2t^{2}}{s_{r}}\Big).$
Since $\mathbb{E}[T_{r}\mid\mathcal{G}]=\sum_{j\in S_{r}}\mathbb{E}[Y_{j}\mid\mathcal{G}]\geq\sum_{j\in S_{r}}\gamma_{j}$, choose
$c_{m}:=\sqrt{\frac{1}{2}\log\frac{m}{\delta_{\mathrm{cnt}}}}\leq\sqrt{\frac{1}{2}\log\frac{3m}{\delta}},\quad t_{r}:=c_{m}\sqrt{s_{r}}.$
Then $\mathbb{P}(T_{r}\geq\sum_{j\in S_{r}}\gamma_{j}-c_{m}\sqrt{s_{r}}\mid\mathcal{G})\geq 1-\delta_{\mathrm{cnt}}/m$. A union bound over $r=1,\dots,m$ yields an event $\mathcal{E}_{\mathrm{cnt}}$ with $\mathbb{P}(\mathcal{E}_{\mathrm{cnt}})\geq 1-\delta_{\mathrm{cnt}}$ on which, for all $r$,
$T_{r}\geq\sum_{j\in S_{r}}\gamma_{j}-c_{m}\sqrt{s_{r}}.$ (B.3)
Combining (B.2) and (B.3) gives, for all $r$,
$T_{r}\geq s_{r}\big(\bar{\gamma}_{m}-\varepsilon_{\mathrm{al}}(\alpha_{r},\delta)\big)-c_{m}\sqrt{s_{r}}$
on the event $\mathcal{E}_{\mathrm{al}}\cap\mathcal{E}_{\mathrm{cnt}}$. Using $s_{r}\geq r=m\alpha_{r}$ and $c_{m}/\sqrt{m}=b_{m}(\delta)$ yields
$U_{r}\geq T_{r}\geq m\Big(\big(\bar{\gamma}_{m}-\varepsilon_{\mathrm{al}}(\alpha_{r},\delta)\big)\alpha_{r}-b_{m}(\delta)\sqrt{\alpha_{r}}\Big).$
Equivalently, at least $m\alpha_{\mathrm{eff}}(\alpha_{r})$ of the $\hat{\Delta}_{j}$’s exceed $\bar{V}_{m}(1-\alpha_{r})$, hence
$\hat{V}_{m}\big(1-\alpha_{\mathrm{eff}}(\alpha_{r})\big)\geq\bar{V}_{m}(1-\alpha_{r}),\quad r=1,\dots,m.$ (B.4)
Finally, we work on $\mathcal{E}_{\mathrm{DKW}}\cap\mathcal{E}_{\mathrm{al}}\cap\mathcal{E}_{\mathrm{cnt}}$, which has probability at least $1-\delta$ by the union bound. Fix any $\alpha\in(0,1)$ and let $\alpha_{+}:=\lceil m\alpha\rceil/m$. By discretization, $\alpha_{\mathrm{eff}}(\alpha)=\alpha_{\mathrm{eff}}(\alpha_{+})$. Applying (B.4) at $r=\lceil m\alpha\rceil$ yields
$\hat{V}_{m}\big(1-\alpha_{\mathrm{eff}}(\alpha)\big)=\hat{V}_{m}\big(1-\alpha_{\mathrm{eff}}(\alpha_{+})\big)\geq\bar{V}_{m}(1-\alpha_{+}).$
Applying (B.1) with $u=1-\alpha_{+}$ gives
$\mathbb{P}_{\psi\sim\Psi}\big(\Delta(\psi)\leq\bar{V}_{m}(1-\alpha_{+})\big|\mathcal{D}\big)$ $\geq 1-\alpha_{+}-\varepsilon_{m}(\delta)$
$\geq 1-\alpha-\varepsilon_{m}(\delta)-\frac{1}{m}.$
Monotonicity in the threshold implies (3.6), hence shown. ∎
Corollary B.1.
Adopt the setting of Theorem (3.1) and set $\gamma_{j}=\gamma$ for all $j\in[m]$. Then, for any $\alpha\in(0,1)$, with probability at least $1-\eta$ over $\mathcal{D}$, we have
$\mathbb{P}_{\psi\sim\Psi}\Big(\Delta_{\psi}^{(k)}\leq\hat{V}_{m}\!\left(1-\gamma\alpha\right)\Big|\mathcal{D}\Big)\geq 1-\alpha-\frac{\varepsilon_{\gamma}(\alpha,m,\eta)}{\sqrt{m}},$ (B.5)
where the remainder $\varepsilon_{\gamma}$ is $O\big(\sqrt{(\log m)/m}\big)$ as $m\to\infty$.
B.2 Proof for Corollary B.1:
@@
One method is to directly adopt Theorem 3.1’s proof and readjust the remainder term. Here we adopt a slightly alternative method. We first consider a fixed level of confidence $\bar{\alpha}$, then extend to $\bar{\alpha}$ holding uniformly across $(0,1)$. Throughout the proof, $\alpha\in(0,1)$ denotes a generic quantile index (a function argument) and is distinct from the target coverage level $\bar{\alpha}$.
As a setup, we work conditionally on $\mathcal{G}_{j}:=\sigma(\psi_{j},\hat{q}_{j},n_{j}),\mathcal{G}:=\sigma\big(\{\mathcal{G}_{j}\}_{j=1}^{m}\big).$ Here $\mathcal{G}_{j}$ is the $\sigma$-field collecting all “human-side” information for scenario $j$. Thus $\mathcal{G}:=\sigma(\{\mathcal{G}_{j}\}_{j=1}^{m})$ represents the joint information from all scenarios that we treat as fixed when analyzing the remaining randomness coming from the estimation noise in $\hat{p}_{j}$. For brevity, we abuse notation and write $p(\psi_{j}),q(\psi_{j})$ as $p_{j},q_{j}$, and similarly $\hat{p},\hat{q}$. We also ignore the superscript $k$ to simplify the proof, and discussions on the dependence on $k$ can be found in Section 3. In addition, for any quantities $\{\Delta_{j}\}_{j=1}^{m}$, we denote the sorted version as $\{\Delta_{(i)}\}_{i=1}^{m}$, ie. $\Delta_{(1)}\leq\cdots\leq\Delta_{(m)}.$ For any sequence $\{\Delta_{j}\}_{j=1}^{m}$, let $\Delta_{(1)}\leq\cdots\leq\Delta_{(m)}$ denote its order statistics.
@@
However, for the above argument to hold, we would need to know $p_{j}$, instead we only have the estimated $\hat{p}_{j}$. Therefore, we instead construct a pseudo-gaps using confidence sets with coverage level $\gamma$, and define the two gap terms as
@@
where $\mathcal{C}_{j}\!\left(\hat{p}_{j},\gamma\right)\subset\Theta$ are data-driven confidence sets satisfying $\mathbb{P}(p_{j}\in\mathcal{C}_{j}\!\left(\hat{p}_{j},\gamma\right)\big|\psi_{j},n_{j})\geq\gamma.$
By Assumption 3.2 and the compactness of the confidence set $C_{j}(\hat{p}_{j},\gamma)\subset\Theta$, Berge’s maximum theorem guarantees that the supremum in the definition of $\hat{\Delta}_{j}$ is attained, hence $\hat{\Delta}_{j}$ is well defined. Therefore, by the coverage property of $C_{j}$, we have $\mathbb{P}(\hat{\Delta}_{j}\geq{\Delta}_{j}|\mathcal{G}_{j}\cup\sigma(\gamma))\geq\gamma,$ and we get $\forall j$
$\mathbb{P}(\hat{\Delta}_{j}\geq{\Delta}_{j}|\mathcal{G}_{j})\geq\gamma.$ (B.7)
@@
$\mathbb{P}\big(\hat{\Delta}_{j}\geq\Delta_{j}\big)=\mathbb{\mathbb{E}}\big[\mathbb{P}(\hat{\Delta}_{j}\geq\Delta_{j}\big|\mathcal{G}_{j})\big]\geq\gamma.$
Moreover, by the above discussion, conditional on $\mathcal{G}$, the indicators $Y_{j}=\mathbf{1}\{\hat{\Delta}_{j}\geq\Delta_{j}\}$ are independent across $j$. We will use $\{\hat{\Delta}_{j}\}_{j=1}^{m}$ to create an upper bound of $\bar{V}_{m}(1-\alpha)$, which along side (B.2) will give us our desired envelope. Intuitively, we want to find a larger quantile of $\hat{\Delta}_{j}$ and with (B.7), we can claim an upper bound of $\bar{V}_{m}(1-\alpha)$ with high probability.
@@
We next define $Y_{j}=\mathbb{I}(\hat{\Delta}_{j}\geq\Delta_{j})$. By (B.7): $\mathbb{P}(Y_{j}=1|\mathcal{\mathcal{G}})\geq\gamma$. Next, define $T_{\alpha}=\{j\in S_{\alpha}:\hat{\Delta}_{j}\geq\Delta_{j}\}=\{j\in S_{\alpha}:Y_{j}=1\}=\{j:\hat{\Delta}_{j}\geq\Delta_{j}\geq\bar{V}_{m}(1-\alpha)\}$. First, we calculate $\mathbb{E}[T_{\alpha}]$:
$\mathbb{E}[T_{\alpha}|\mathcal{G}]=\mathbb{E}\Big[\sum_{j\in S_{\alpha}}Y_{j}|\mathcal{G}\Big]=\sum_{j\in S_{\alpha}}\mathbb{P}(Y_{j}|\mathcal{G})\geq\gamma s.$
Lemma A.2 implies that for any $\delta\in[0,\gamma s]$, we have:
$\mathbb{P}\Big(|T_{\alpha}|\leq\gamma s-t|\mathcal{G}\Big)$ $\leq\mathbb{P}\Big(|T_{\alpha}|\leq\mathbb{E}[T_{\alpha}]-t|\mathcal{G}\Big)$
@@
where we applied $\mathbb{E}[T_{\alpha}|\mathcal{G}]\geq\gamma s$ in the first inequality. By setting $t=c\sqrt{s}$:
$\mathbb{P}\Big(|T_{\alpha}|\leq\gamma s-t|\mathcal{G}\Big)\leq\exp(-2c^{2})$ (B.8)
@@
$|U_{\alpha}|\geq|T_{\alpha}|\geq\gamma s-c\sqrt{s},$
which implies at least $\gamma s-c\sqrt{s}$ of the $\hat{\Delta}_{j}$’s are larger than $\bar{V}_{m}(1-\alpha)$ with high probability.
We now analyze what coverage guarantee can we get for the inner probability via order statistics for any $\alpha$. Set $N:=\big\lfloor\gamma s-c\sqrt{s}\big\rfloor$ and define $\hat{V}_{m}(\alpha):=\inf\{t:\hat{F}_{m}(t)\geq\alpha\}=\Delta_{(\lceil m\alpha\rceil)}$, where $\hat{F}(x)=\frac{1}{m}\sum_{j=1}^{m}\mathbb{I}(\hat{\Delta}_{j}\leq x)$. If at least $N$ sample values exceed $\bar{V}_{m}(1-\alpha)$, then by order-statistics calculus (Lemma A.7)
@@
$\gamma m\alpha_{\mathrm{eff}}-c\sqrt{m\alpha_{\mathrm{eff}}}-1\geq m\alpha.$
Define $\alpha_{\mathrm{eff}}(\alpha,c,m):=\inf\{x\in(0,1):\gamma x-c\sqrt{\tfrac{x}{m}}-\tfrac{1}{m}\geq\alpha\}$. Writing $y=\sqrt{x}$, this is equivalent to $\gamma y^{2}-\tfrac{c}{\sqrt{m}}y-\big(\tfrac{1}{m}+\alpha\big)\geq 0$, so the minimal admissible $y$ is
$y^{*}$ $=\frac{c/\sqrt{m}+\sqrt{c^{2}/m+4\gamma\alpha+4\gamma/m}}{2\gamma},$
@@
$\text{then}\qquad\mathbb{P}(\mathcal{E}_{\alpha})\geq 1-e^{-2c^{2}}.$
Next, we apply (B.2) at level $1-\alpha_{\mathrm{eff}}(\bar{\alpha},c,m)$:
$\mathbb{P}_{\psi\sim\Psi}\Big(\Delta_{\psi}\leq\bar{V}_{m}(1-\alpha_{\mathrm{eff}})\Big|\mathcal{D}\Big)$
$\geq 1-\alpha_{\mathrm{eff}}(\bar{\alpha},c,m)-\varepsilon_{m}(\delta).$
@@
$\mathbb{P}_{\psi\sim\Psi}\big(\Delta_{\psi}\leq\hat{V}_{m}(1-\bar{\alpha})\big|\mathcal{D}\big)\geq 1-\alpha_{\mathrm{eff}}(\bar{\alpha},c,m)-\varepsilon_{m}(\delta),$ (B.10)
@@
$\alpha_{\mathrm{eff}}(\bar{\alpha},$ $c,m)=\frac{\big(\tfrac{c}{\sqrt{m}}+\sqrt{\tfrac{c^{2}}{m}+4\gamma\bar{\alpha}+\tfrac{4\gamma}{m}}\big)^{2}}{4\gamma^{2}}=$
$\frac{\bar{\alpha}}{\gamma}+\frac{c}{\gamma\sqrt{m}}\sqrt{4\gamma\bar{\alpha}+\tfrac{c^{2}+4\gamma}{m}}+\frac{c^{2}}{2\gamma^{2}m}+\frac{1}{\gamma m}.$
Therefore, as $m\to\infty$, $\alpha_{\mathrm{eff}}(\alpha,c,m)=\tfrac{\bar{\alpha}}{\gamma}+c\sqrt{4\bar{\alpha}/\gamma m}+O(m^{-1})$.
@@
By DKW, $\mathbb{P}(\mathcal{E}_{\mathrm{DKW}})\geq 1-\delta$, and by the fixed-level argument, $\mathbb{P}(\mathcal{E}_{r})\geq 1-e^{-2c^{2}}$ for each $r$. Hence, by a union bound,
$\mathbb{P}\Big(\mathcal{E}_{\mathrm{DKW}}\cap\bigcap_{r=1}^{m}\mathcal{E}_{r}\Big)\geq 1-\delta-me^{-2c^{2}}.$
@@
Applying (B.10) at level $\alpha_{+}$ yields
@@
$\geq 1$ $-\frac{\alpha_{+}}{\gamma}-\frac{c}{\gamma\sqrt{m}}\sqrt{4\gamma\alpha_{+}+\tfrac{c^{2}+4\gamma}{m}}$
$-\frac{c^{2}}{2\gamma^{2}m}-\frac{1}{\gamma m}-\varepsilon_{m}(\delta).$
@@
$\frac{\alpha_{+}}{\gamma}\leq\frac{\alpha}{\gamma}+\frac{1}{\gamma m},\quad\sqrt{4\gamma\alpha_{+}+\tfrac{c^{2}+4\gamma}{m}}\leq\sqrt{4\gamma\alpha+\tfrac{c^{2}+8\gamma}{m}}.$
Therefore, with probability at least $1-\delta-me^{-2c^{2}}$ over $\mathcal{D}$, the guarantee (B.5) (at $\alpha$ replaced by $\alpha_{+}$) holds uniformly for all $\alpha\in(0,1)$. The form in the main theorem is a simplification with respect to $c$ and $\delta$ while taking $\gamma=\tfrac{1}{2}$.
B.3 Corollary for simulator budget: